A fast-food company is preparing beef patties for their burgers. The lower specification limit for the burger is 43.6 grams and the upper specification limit is 46.4 grams. The standard deviation is 0.5 grams and the mean is 45 grams.

Based on the information given above, please answer the following 2 questions.

1.) The process capability index for this beef patty production process

p=       .

Keep TWO decimals in your answer.

2.) Please calculate the PPM for this beef patty production process.

Round your answer to the nearest integer.

Note: you need to calculate the defect probability first to calculate PPM.
Answer #1

1. Here USL is 46.4 and LSL is 43.6 and the Standard deviation is 0.5, and the mean is 45.

Cpu= (46.4 - 45)/ (3 * 0.5) = 1.4/1.5 = 0.93

Cpl = (45 - 43.6)/ (3 * 0.5) = 1.4/1.5 = 0.93

Cpk = Minimum of [ Cpu, Cpl]

That is Cpk = minimum of [0.93, 0.93]

The process capability index Cpk is 0.93